JOINT DIFFERENTIAL GEOMETRY/PDE & RAINWATER SEMINAR

Wednesday, November 10, 2004 Padelford C-36 3:50 pm (Non)regularity of projections of measures invariant under geodesic flow

Maarit Järvenpää (University of Jyvaskyla, Finland)

We show that, unlike in the 2-dimensional case, the Hausdorff dimension of a measure invariant under the geodesic flow is not necessarily preserved under the projection from the unit tangent bundle onto the base manifold if the base manifold is at least 3-dimensional. In the 2-dimensional case we reprove the preservation theorem due to Ledrappier and Lindenstrauss using the general projection formalism of Peres and Schlag. The novelty of our proof is that it illustrates the reason behind the failure of the preservation in higher dimensional case. Finally, we show that the projected measure has fractional derivatives of order γ for all γ < (α − 2)/2 provided that the invariant measure has finite α-energy for some α > 2 and the base manifold has dimension 2.

For more information about this seminar, visit the DG/PDE Seminar Web page (from the Math Department home page, www.math.washington.edu, follow the link Seminars, Colloquia, and Conferences).

The University of Washington is committed to providing access, equal opportunity and reasonable accommodation in its services, programs, activities, education and employment for individuals with disabilities. To request disability accommodation contact the Disability Services Office at least ten days in advance at: 206-543-6450/V, 206-543-6452/TTY, 206-685-7264 (FAX), or dso@u.washington.edu.